OpenStudy (anonymous):

can you solve this for me? inverse laplace transform: f(s) = 10/s(s+2)(s+3)^2

hartnn (hartnn):

do you know partial fractions ?

OpenStudy (anonymous):

A-Level Pure Math 3

OpenStudy (anonymous):

@hartnn I know a little bit

OpenStudy (anonymous):

@edwinls9 what do you mean?

hartnn (hartnn):

\(\Large \dfrac{10}{s(s+2)(s+3)^2}=\dfrac{A}{s}+\dfrac{B}{s+2}+\dfrac{C}{s+3}+\dfrac{D}{(s+3)^2}\) can you find A,B,C,D ??

hartnn (hartnn):

10 = A (s+2)(s+3)^2 + Bs (s+3)^2 + Cs(s+2)(s+3)+Ds(s+2) That is true for every value of s put s=0, 10 = A(2)(3)^2 =.... Find A put s = -2 10 = B(-2)(1)^2 =.... Find B put s= -3 10 = D(-3)(-1)=.... Find D put s= -1 10 = A(1)(2)^2 + B(-1)(2)^2 +C(-1)(1)(2)+D(-1)(1) Find C @fariza

hartnn (hartnn):

did you try finding A,B,C, D ?? what values you got ?

OpenStudy (anonymous):

I'm not finding it yet cuz I don't know partial fractions :/

undeadknight26 (undeadknight26):

http://www.mathsisfun.com/algebra/partial-fractions.html

OpenStudy (anonymous):

@undeadknight26 I'm studying this note and trying to understand .

hartnn (hartnn):

ok, are you clear till here, \(\Large \dfrac{10}{s(s+2)(s+3)^2}=\dfrac{A}{s}+\dfrac{B}{s+2}+\dfrac{C}{s+3}+\dfrac{D}{(s+3)^2}\) ?

OpenStudy (anonymous):

yes @hartnn

hartnn (hartnn):

now do you know how to make a common denominator on right side ?

OpenStudy (anonymous):

no

hartnn (hartnn):

ok, lets take example of only 2 fraction, say, a/b + c/d how would you simplify that ? nake a common denominator

OpenStudy (anonymous):

ad+cb/bd true?

hartnn (hartnn):

yes so, how about 3 fractions ? a/b +c/d + e/f = ... ?

hartnn (hartnn):

common denominator will be bdf right ? what about numerator ?

OpenStudy (anonymous):

I don't know

hartnn (hartnn):

okk, lets go other way round, \(\Large \dfrac{10}{s(s+2)(s+3)^2}=\dfrac{A}{s}+\dfrac{B}{s+2}+\dfrac{C}{s+3}+\dfrac{D}{(s+3)^2}\) now multiply EVERY term by s (s+2) (s+3)^2 what do u get ?

OpenStudy (anonymous):

do I have to expand it?

hartnn (hartnn):

just multiply each term by s(s+2)(s+3)^2 do not expand

OpenStudy (anonymous):

what answer do u get?

hartnn (hartnn):

\(\Large \dfrac{10}{s(s+2)(s+3)^2} \times s(s+2)(s+3)^2 \\ \Large =\dfrac{A}{s} \times s(s+2)(s+3)^2+\dfrac{B}{s+2}\times s(s+2)(s+3)^2 \\ \Large +\dfrac{C}{s+3}\times s(s+2)(s+3)^2+\dfrac{D}{(s+3)^2}\times s(s+2)(s+3)^2\) simplify each term, cancel whatever gets cancelled from numerator and denominator. example, on left side, only 10 remains

hartnn (hartnn):

let me help you get started, \(\Large \dfrac{10}{ \cancel {s(s+2)(s+3)^2}} \times \cancel {s(s+2)(s+3)^2 }\\ \Large =\dfrac{A}{\cancel s} \times \cancel s(s+2)(s+3)^2+\dfrac{B}{\cancel {s+2}}\times s\cancel {(s+2)}(s+3)^2 \\ \Large +\dfrac{C}{\cancel {s+3}}\times s(s+2)(s+3)^{\cancel 2}+\dfrac{D}{\cancel {(s+3)^2}}\times s(s+2)\cancel{(s+3)^2}\)

OpenStudy (anonymous):

so what the next step? I'm trying to understand this

hartnn (hartnn):

did you get how those terms are getting cancelled ?

hartnn (hartnn):

next step : \(10 = A (s+2)(s+3)^2 + Bs (s+3)^2 + Cs(s+2)(s+3)+Ds(s+2) \)

OpenStudy (anonymous):

ok then?

hartnn (hartnn):

you got that ? because after that, things are easy. that is true for all values of s so, put s=0 in that and tell me what u get?

OpenStudy (anonymous):

why s=0?

hartnn (hartnn):

because we can choose any value of s if we choose s = 0 then B,C,D will vanish and we will be able to find A :)

OpenStudy (anonymous):

the 's' of 'Bs,Cs and Ds must be 0 too?

hartnn (hartnn):

yes

hartnn (hartnn):

could you get A ?

OpenStudy (anonymous):

18?

hartnn (hartnn):

10 = A (18) you got that, isn't it ?

OpenStudy (anonymous):

Yaa I got it . so the next step is?

hartnn (hartnn):

10 = A (s+2)(s+3)^2 + Bs (s+3)^2 + Cs(s+2)(s+3)+Ds(s+2) That is true for every value of s put s=0, 10 = A(2)(3)^2 =.... Find A put s = -2 10 = B(-2)(1)^2 =.... Find B put s= -3 10 = D(-3)(-1)=.... Find D put s= -1 10 = A(1)(2)^2 + B(-1)(2)^2 +C(-1)(1)(2)+D(-1)(1) Find C

hartnn (hartnn):

find B,C,D

OpenStudy (anonymous):

why the value of s is different?

hartnn (hartnn):

s can take any value wee need to choose we choose value of s such that every other constant gets eliminated

OpenStudy (anonymous):

okay :]

OpenStudy (anonymous):

Usually it's a requirement to have Calc II on board before Differential Equations..

OpenStudy (anonymous):

@FutureMathProfessor means?